• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 08:01:08 on June14,2010

    Entry number 326450

    keyword=June 14 Owl Shift Summary

    00:00----Shift starts with transverse polarization measurements. KK called about the run plan. Dustin and Jon will be here.

    00:44----Hall is in controlled access. Jon and Megan are going to hall to blind the PMTs. High voltages of all the detectors are off. Heater is ON and cooler is OFF.

    02:11----Hall is in pebam permit.

    03:58----Dustin and Jon are in the hall to unblind the PMTs.

    04:42----Hall is in beam permit.

    05:32----Half wave plate is out and run the flipper.

    05:50----Blind test ends.